# revtex4-2Repair the float

### v1

@@ III.1 Baseline algorithms
To understand how standard algorithms fare with such a complex posterior distribution, we first tested a naive random walk MCMC (RW) algorithm (details in Appendix C) on a small synthetic graph with known ground truth embedding coordinates (shown in Fig. 1). We quantified our results using the autocorrelation, the effective number of samples $n_{\mathrm{eff}}$, and $\hat{R}$—a measure of mixing (see Appendix D for details). Further, we tested convergence by initializing the algorithm (i) at the ground truth and (ii) with a simple initialization strategy that can be applied to observational data. Figure 3 illustrates that although the RW has a high autocorrelation, it very slowly reaches the typical set—the part of the distribution that holds a significant probability mass and where good algorithms should naturally spend most of its time  [63].
@@
For our second test, we wrote the model in Stan, a probabilistic programming language that implements a state-of-the-art dynamic Hamiltonian Monte Carlo (HMC) algorithm [64] for sampling from arbitrary Bayesian models. Dynamic HMC runs slowly because each sampling step involves integrating a large system of differential equations. Nonetheless, the promise of high-quality samples makes this option worth exploring.
@@
Unfortunately, the barriers shown in Fig 2 are problematic for a purely gradient-based method such as HMC, and Stan indeed reports gradient divergences, which signal incorrect sampling. This issue can be partially mitigated by using an approximate likelihood (described in Appendix B) that smoothens these barriers at the cost of a slightly distorted posterior. However, even with this modification, the algorithm is unable to explore the sampling space sufficiently fast (we reach the maximum tree depth easily [64]).
@@
Stan’s default parameters can be modified to allow for better exploration (increased tree depth), but even then, a critical problem remains and seems insurmountable: after a large number of iterations, the sampled distribution differs depending on the initialization, a signal that the mixing is poor; see Fig. 3. Initializations at the ground truth yield high-quality samples, but random initializations do not converge to the typical set even after 200,000 iterations. This mixing issue is likely caused by the complicated and non-convex geometry of the posterior distribution, which motivates the modifications we propose next.
III.2 Cluster transformation MCMC
@@
We can improve upon dynamic HMC and RW by noticing that the connection probability is a decreasing function of angular separation, such that groups of vertices at a large enough distance from one another are essentially independent. As a result, each region is akin to a smaller and simpler embedding problem, which is the basis for the approach of Wang et al. [41, 62]. (This is also the basis for the idea of renormalization in the $\mathbb{H}^{2}$ hyperbolic model in which groups of vertices are coarse-grained while partially preserving the structural integrity of the graph [14, 1].) Recalling that the likelihood is invariant to reflections and rotations, we propose to apply these transformations at the level of groups or clusters of vertices.
@@
The first step for these new transformations is to partition vertices into roughly independent clusters given the angular positions. This is done by grouping vertices with an angular separation below some randomly sampled threshold in clusters. We use this naive partitioning scheme instead of an optimization procedure [65] for technical reasons discussed in Appendix C.2.
@@
Our proposed algorithm, BIGUE, combines these random cluster transformations (CT) with the random walk (RW) baseline. Cluster transformations explore the mesoscale structure of the embedding, while the random walk fine-tunes the embedding coordinates, including the parameters $\{\kappa_{u}\}$. Each sampling step is either a randomly selected CT or a short RW, and we calibrate the transition probabilities to the posterior distribution using the Metropolis-Hastings algorithm; see Appendix C for the technical details. (One could also combine random cluster transformations with HMC, but the high computational cost of HMC turns out to outweigh its benefits. A naive random walk provides cheaper and sufficient fine-tuning in our experiments.) We deal with the symmetries and automorphisms described in Sec. II.2 with a post-processing alignment step, in which we minimize the sum of angular separations to some arbitrary reference embedding over the automorphisms found by Nauty [66], the two possible reflections (reflection or no reflection) and the rotations $[0,2\pi)$.
@@
Figure 3 shows that BIGUE has a lower autocorrelation and reaches the typical set much faster than a pure RW algorithm. The fluctuations at equilibrium are also larger, which suggests that it explores the posterior more efficiently. Furthermore, the maximum potential reduction factor $\hat{R}_{\text{max}}$ of BIGUE is lower than that of a random walk, suggesting that the algorithm has better mixing.
When comparing HMC to the complete Markov Chain generated with BIGUE, we note that autocorrelation decays far more slowly. Furthermore, BIGUE yields $\hat{R}$ values above the recommended $1.01$ threshold [67] (as do all the other algorithms). A direct fix is to thin the BIGUE chain, that is, only use every $k$-th sample point of the chain. Guided by Fig. 3a, we set $k=10,000$ and obtain a $\hat{R}_{\text{max}}<1.01$ in, with the chain shown in Fig. 3c.
We note that Fig. 3 highlights one unintuitive aspect of continuous random variables: The neighborhood of the likelihood maximum accounts for a small fraction of the probability mass because probabilities are obtained through integration with respect to the Lebesgue measure. While the log-likelihood maximum (found by Mercator [33]) has a high probability density, its surroundings cover a larger volume (the typical set) and hold a much larger probability mass. Hence, our MCMC sample concentrates away from the maximum likelihood, as expected. (That said, we note that cluster transformations could easily be added to maximization routines if a point estimate was the goal, something we leave for future work.)
IV Uncertain embeddings
IV.1 Embedding and network properties
@@
Figure 4: Posterior estimates and posterior predictive distribution for the synthetic graph used in Fig. 3. (a) Angular coordinates $\{\theta_{u}\}$. (b) Parameters $\{\kappa_{u}\}$. (c) Inverse temperature $\beta$. (d) Density. (e) Clustering (the transitivity). (f) Greedy routing success rate. (g) Global hierarchy level [68]. (h) AUC of link prediction with embeddings coordinates and the $\mathbb{S}^{1}$ likelihood. In each case, $500$ embeddings were sampled from four independent chains. In panels (c-h), the vertical dotted lines show point estimates computed with the Mercator and ground truth embeddings. In panels (a,b), points denote the median, and the error bars cover the first to third quartile. In panels (d,e), we generated graph samples for Mercator and the ground truth by conditioning the model’s likelihood on point estimates of the embedding.
@@
Figure 5: Multimodality and non-normality of the posterior distribution, shown with the marginal posterior distributions of pairs of angular coordinates in the embedding of (a) the synthetic graph in Fig. 3 (b) Zachary’s karate club [69] (c) a synthetic graph containing a vertex whose position is drawn from a mixture. Each point is obtained with the thinned BIGUE algorithm, and chains comprise $500$ random embeddings. The black lines delineate the density of a normal kernel density estimation of the marginal. To account for the periodicity, the kernel density is summed for rotations ±$\pm 2\pi$ in both axes.
@@
Figure 6: Comparison of network embeddings obtained from a maximum likelihood estimator (Mercator) and the posterior of a Bayesian model (BIGUE) for various empirical networks: Zachary’s karate club [69], cortical connectivity of the macaque [70], street gangs in Montreal [71], dutch literary critics [72], zebra social interactions [73], connections in greek terrorist group [74], kagaroo dominance relationships [74] and New Guinea tribes friendships. Vertices outside the largest connected components were removed in order to run Mercator. The horizontal bars show 50% highest density intervals, and points within are the median. Table 1 reports the values for this figure.
@@
A proper quantification of the uncertainty inherent to the data is the primary reason for using a Bayesian method such as BIGUE. Figure 4 shows what such an analysis might look like, using credible intervals for the embedding coordinates and their transformations, computed with samples drawn from the chains depicted in Figure 3b.
@@
The first to note is that BIGUE recovers the ground-truth coordinates, and so does Mercator. Posterior sampling reveals some leeway in the coordinates compatible with the data $G$, but they are firmly centered around the ground-truth values (Fig. 4a). As expected, for most vertices $u$, the $\kappa_{u}$ parameters concentrate on the vertices’ original degrees (Fig. 4b), but there are discrepancies. The relationship $\mathbb{E}[K(\kappa_{u})]=\deg(u)$ is obtained by integrating the vertex density on the circle, which assumes an infinite number of vertices. This is why, for a high-degree vertex $u$ in a small graph, $\kappa_{u}$ is larger to compensate for the fact that there is a finite number of vertices. Figure 4c also shows that the posterior covers the original $\beta$ value, although it is in a somewhat low probability region. The posterior distribution agrees with Mercator and places a higher likelihood than the ground truth on $\beta$ being a bit under $2$—this is due to random fluctuations in the instantiation of $G$ conditioned on the ground truth coordinates.
@@
Next, we turn to the properties of the graph. The edge probabilities of good embeddings should yield graphs with properties similar to those of the original graph. As a result, if the embedding is accurate, we expect that the graphs generated with estimated coordinates will be similar to the original graph. We first verify this with Mercator, an optimization algorithm that provides a single embedding and, thus, a single random graph ensemble; the resulting distribution of graph properties is centered around the observed graph in Fig. 4d,e. With BIGUE, each sample point corresponds to a random graph using the likelihood, and consequently, we can marginalize the graph properties over a much larger parameter space. (Formally, we can draw from the posterior predictive distribution of the model.). The resulting distributions are wider and now capture parameter uncertainty; see Fig. 4d,e again.
@@
The embeddings also capture less obvious properties. For example, hyperbolic embeddings can play a role in finding effective and efficient greedy routing on the network. In the greedy routing algorithm, hyperbolic coordinates act as addresses, and one attempts to reach the target $v$ from a source $u$ by repeatedly following the edge that leads to the neighbor closest to the destination $u$ (in hyperbolic space). The main issue with this algorithm is that paths can sometimes devolve into “greedy loops” that lead nowhere. The greedy success rate evaluates the extent to which this is an issue as the proportion of greedy routes that successfully reach their destination. Our posterior samples allow us to compute error bars for this property of the graph and embedding, as shown in Fig. 4f. In this particular synthetic graph, many plausible embeddings lead to a greedy routing comparable to Mercator, but the majority is less reliable.
@@
The hierarchy of graphs is another example of complex properties captured by hyperbolic embeddings. Garcia et al. [68] introduced the global hierarchy level to quantify this, using a function of the angular distance between neighboring nodes with coordinates in the outer and inner parts of the embedding. Figure 4g illustrates how, for the synthetic graph analyzed, most plausible embeddings have smaller global hierarchy levels than both the ground truth and the Mercator embedding.
Finally, embeddings for the $\mathbb{S}^{1}$ model can be viewed as classifiers for edges since the likelihood maps potential edges (binary classes) to edge probabilities. If the embedding is a good representation of the graph, edges should have a high probability, while non-edges should have a low probability. As a simple test of link prediction, we used the embedding positions and the likelihood (II.1) to assign probabilities to pairs of nodes. We then quantified the quality of this classifier with the area under the receiver operating characteristic curve (AUC), with a score of $1$ indicating perfect edge prediction. Figure 4h reveals that Mercator’s embedding yields the best classifier, while BIGUE’s embeddings are as useful as the ground-truth coordinates. This suggests that BIGUE reports more typical embeddings of the model while Mercator overfits the coordinates to reproduce exactly the same graph.
IV.2 Multimodality

### v2

@@ II.2.1 Baseline algorithms
We first tested a naive random walk MCMC (RW) algorithm (details in Supplementary Methods 1) on a small synthetic graph with known ground truth embedding coordinates (shown in Fig. 1). We quantified our results using the normalized autocovariance, the effective number of samples $S_{\mathrm{eff}}$, and $\hat{R}$—a measure of mixing (see Supplementary Methods 2 for details). Furthermore, we tested convergence by initializing the algorithm (i) at the ground truth and (ii) with a simple initialization strategy that can be applied to observational data. Figure 3 illustrates that the RW has a high normalized autocovariance and it very slowly reaches the typical set—the part of the distribution that holds a significant probability mass and where good algorithms should naturally spend most of their time  [73].
@@
For our second test, we wrote the model in Stan, a probabilistic programming language that implements a dynamic Hamiltonian Monte Carlo (HMC) algorithm [74] for sampling from arbitrary Bayesian models. Dynamic HMC runs slowly because each sampling iteration involves integrating a large system of differential equations. Nonetheless, the promise of high-quality samples makes this option worth exploring.
@@
Unfortunately, the barriers shown in Fig 2 are problematic for a purely gradient-based method such as HMC, and Stan indeed reports gradient divergences, which signal incorrect sampling. This issue can be partially mitigated by using an approximate likelihood (described in Supplementary Note 3) that smoothens these barriers at the cost of a slightly distorted posterior. However, even with this modification, the algorithm is unable to explore the sampling space sufficiently fast (we reach the maximum tree depth easily [74]).
@@
Stan’s default parameters can be modified to allow for better exploration (increased tree depth), but even then, a critical problem remains and seems insurmountable: after a large number of iterations, the sampled distribution differs depending on the initialization, a signal that the mixing is poor; see Fig. 3. Initializations at the ground truth yield samples of higher quality than the RW, but random initializations do not converge to the typical set even after $200,000$ iterations. This mixing issue is likely caused by the complicated and non-convex geometry of the posterior distribution, which motivates the modifications we propose next.
Figure 4: Summary of the cluster transformations used in BIGUE. The flip transformation targets the reflection symmetries, and the exchange and translate transformations target the rotation symmetries.
II.2.2 Cluster transformation MCMC
@@
We can improve upon dynamic HMC and RW by noticing that the connection probability is a decreasing function of the angular separation, such that groups of vertices at a large enough distance from one another are essentially independent. As a result, each region is akin to a smaller and simpler embedding problem, which is the basis for the approach of Wang et al. [44, 71]. (This is also the basis for the idea of renormalization in the $\mathbb{H}^{2}$ hyperbolic model in which groups of vertices are coarse-grained while partially preserving the structural integrity of the graph [17, 1].) Recalling that the likelihood is invariant to reflections and rotations, we propose to apply these transformations at the level of groups or clusters of vertices; c.f. Fig. 4.
@@
The first step for these transformations is to partition vertices into roughly independent clusters given the angular positions. This is done by grouping vertices with an angular separation below some randomly sampled threshold in clusters. We use this naive partitioning scheme instead of an optimization procedure (e.g. modularity maximization [75]) for technical reasons discussed in Supplementary Methods 1.
@@
Our proposed algorithm, BIGUE, combines these random cluster transformations (CT) with the random walk (RW) baseline. Cluster transformations explore the mesoscale structure of the embedding, while the random walk fine-tunes the embedding coordinates, including the parameters $\kappa$. Each sampling iteration is either a randomly selected CT or a RW, and we calibrate the transition probabilities to the posterior distribution using the Metropolis-Hastings algorithm; see Supplementary Methods 1 for the technical details. (One could also combine random cluster transformations with dynamic HMC, but the high computational cost of dynamic HMC turns out to outweigh its benefits. A naive random walk provides cheaper and sufficient fine-tuning in our experiments.)
@@
Figure 3 shows that BIGUE has the lowest autocovariance and reaches the typical set faster the other algorithms. The fluctuations at equilibrium are also larger, which suggests that it explores the posterior more efficiently. Furthermore, BIGUE’s maximum potential reduction factor $\hat{R}_{\text{max}}$ is the lowest, suggesting that it has better mixing. However, the three algorithms yield $\hat{R}$ values above the recommended $1.01$ threshold [76].
A direct fix to high normalized autocovariance and potential reduction factor is to thin the chains, that is, only use every $k$-th sample point of the chain. Guided by Fig. 3a, we set $k=10,000$ for BIGUE and the RW and obtain the chains of panel d. We now see that BIGUE’s chains have $\hat{R}_{\text{max}}<1.01$ while those of the RW do not. Moreover, the fluctuations of the RW are still smaller, indicating, again, that the sampling space is better explored by BIGUE.
From the samples of Fig. 3, we also compute the effective sample size of the chains, which is related to the estimation error in the Markov Chain central limit theorem. The median effective sample sizes $S_{\text{eff}}$ across all parameters were, for BIGUE: $486$, $462$, and $1814$ (in panels b,c, and d respectively). For the random walk MCMC, we obtained medians of $S_{\text{eff}}=248$, $384$, and $1390$, and for dynamic HMC, they were equal to $593$ and $347$. This provides further evidence that cluster transformation benefit the random walk algorithm. Autocovariance, $\hat{R}$ and $S_{\text{eff}}$ and their adaptation to periodic random variables are explained in Supplementary Methods 2.
@@
Figure 5: Posterior estimates and posterior predictive distribution for the synthetic graph used in Fig. 3. (a) Angular coordinates $\theta$. (b) Parameters $\kappa$. (c) Inverse temperature $\beta$. (d) Density. (e) Clustering (the transitivity). (f) Greedy routing success rate. (g) Global hierarchy level [77]. (h) Normalized rank of the removed edges’ existence probability among all the unconnected pairs (see below). In each case, $2000$ total embeddings were sampled from four independent chains. All shades of blue, green and gray display the values for the posterior sample, Mercator and the ground truth, respectively. In panels (a,b), points denote the median, and the error bars cover the interquartile range. In panels (c-g), the vertical dotted lines show point estimates computed with the Mercator and ground truth embeddings. In panels (d,e), we generated graph samples for Mercator and the ground truth by conditioning the model’s likelihood on point estimates of the embedding. In panels (c, f,g), the blue dashed line is the median of the posterior sample. In panel (h), we test the algorithm on a link prediction task [78], and report the rank of removed edges as predicted by the likelihood. Pairs of embedding and graphs of 30 vertices were sampled from the prior. For each graph, $5\%$ of the edges were then randomly removed ($4$ to $5$ edges), and the normalized ranks of removed edges were calculated and binned. (Disconnected graphs were rejected for compatibility with Mercator). The figure reports the median and interquartile range of the normalized rank frequencies calculated across graphs, removals, and samples. The AUC values are shown in Supplementary Note 5.
In light of these simulations, we conclude that BIGUE outperforms both dynamic HMC and the RW in all metrics, and that in general, we expect these algorithms to perform equivalently to BIGUE at best. For this reason, the remainder of the paper uses BIGUE to sample the posterior distribution.
It may seem surprising that BIGUE relies on clusters in the embedding while the vertices are uniformly distributed according to prior density of the generative model. However, the clusters found need not be well separated to be useful in sampling. This is due to the locality of connections: If some pairs of vertices are sufficiently distant, then we can reasonably attempt to separate these vertices as they might not be connected anyway. Conversely, vertices that are closeby in the best embedding will tend to be connected and can thus be moved as a unit instead of one at a time.
We note that Fig. 3 highlights one counterintuitive aspect of continuous random variables [73]: The neighborhood of the likelihood maximum accounts for a small fraction of the probability mass because probabilities are obtained through integration with respect to the Lebesgue measure. While the log-likelihood maximum (found by Mercator) has a high probability density, its surroundings cover a larger volume (the typical set) and hold a much larger probability mass. Hence, our MCMC sample concentrates away from the maximum likelihood, as expected. (That said, we note that cluster transformations could easily be added to maximization routines if a point estimate was the goal, something we leave for future work.)
III Results
Having a reliable algorithm to sample embeddings from the posterior distribution, we can now estimate embeddings and integrals like Eq. (8). This section illustrates how the resulting Bayesian approach compares to maximum likelihood estimation both on a synthetic graph and various empirical networks.
III.1 Embedding and network properties
@@
Figure 6: Multimodality and non-normality of the posterior distribution, shown with the marginal posterior distributions of pairs of angular coordinates in the embedding of (a) the synthetic graph in Fig. 3 (b) Zachary’s karate club [79] (c) a synthetic graph of $30$ vertices with two conflicting embeddings, as described in section III.2 ($\theta_{0}$ has two different ground truth positions). Samples are obtained with the thinned BIGUE algorithm, and chains comprise $500$ random embeddings. The color of each circle indicates from which chain it was sampled. The black lines delineate the density of a normal kernel density estimation of the marginal. The sample averages $\bar{\theta}_{u}$ are circular averages as described in Supplementary Methods 2.
@@
Figure 7: Comparison of network embeddings obtained from a maximum likelihood estimator (Mercator, green cross) and the posterior of a Bayesian model (BIGUE, blue circle) for various empirical networks. These networks are Zachary’s karate club [79], cortical connectivity of the macaque [80], street gangs in Montreal [81], dutch literary critics [82], zebra social interactions [83], connections in greek terrorist group [84], kangaroo dominance relationships [85] and New Guinea tribes friendships [86]. The value for the original network are displayed in gray stars when available. Vertices outside the largest connected component were removed in order to run Mercator. The horizontal bars show the 50% highest density intervals, and points within are the median. Supplementary Table 1 reports the values for this figure.
@@
We first compare our Bayesian model to Mercator on the synthetic graph used in Fig. 3. Figure 5 shows what such an analysis might look like, using credible intervals for the embedding coordinates and their transformations, computed with samples drawn from the chains depicted in Fig. 3d.
@@
We see in Fig. 5a that Mercator’s estimation generally lies within the estimated credible intervals, revealing some leeway in the coordinates compatible with the data $G$. The ground truth value is sometimes not within the interval because the graph is small (see Supplementary Figure 1, where the graph contains $100$ vertices, and the coordinates are recovered almost perfectly). As expected, for most vertices $u$, the $\kappa_{u}$ parameters concentrate on the vertices’ original degrees (Fig. 5b), but there are discrepancies. Note that the relationship $\mathbb{E}[K|\kappa_{u}]=\deg(u)$, mentioned in the Model definition section, is valid assuming an infinite number of vertices. This is why, for a high-degree vertex $u$ in a small graph, $\kappa_{u}$ is larger to compensate for the fact that there is a finite number of vertices. Figure 5c also shows that the posterior covers the original $\beta$ value, although it is in a somewhat low probability region. The posterior distribution agrees with Mercator and places a higher likelihood than the ground truth on $\beta$ being a bit under $2$—this is due to random fluctuations in the instantiation of $G$ conditioned on the ground truth coordinates.
@@
Next, we turn to the properties of the graph. The edge probabilities of good embeddings should yield graphs with properties similar to those of the original graph. As a result, if the embedding is accurate, we expect that the graphs generated with estimated coordinates will be similar to the original graph. We first verify this with the maximum likelihood found with Mercator; the resulting distribution of graph properties is centered around the observed graph in Fig. 5d,e. We note that although Mercator yields a point-wise estimation, different graphs can be sampled from the same embedding, which explains why there are many density and clustering values for Mercator. With BIGUE, each sample point corresponds to a random graph using the likelihood, and consequently, we can marginalize the graph properties over a much larger parameter space. (Formally, we draw from the posterior predictive distribution of the model.) The resulting distributions are wider and now capture parameter uncertainty; see Fig. 5d,e.
@@
The embeddings also capture less obvious properties. For example, hyperbolic embeddings can play a role in finding effective and efficient greedy routing on the network. In the greedy routing algorithm [10], hyperbolic coordinates act as addresses, and one attempts to reach the target $v$ from a source $u$ by repeatedly following the edge that leads to the neighbor closest to the destination $u$ (in hyperbolic space). The main issue with this algorithm is that paths can sometimes devolve into “greedy loops” that lead nowhere. The greedy success rate evaluates the extent to which this is an issue as the proportion of greedy routes that successfully reach their destination. The posterior samples allows us to compute error bars on the greedy success rate, but also to comment the navigability of the ensemble of plausible embeddings. For instance, Fig. 5f shows that, while many embeddings are as navigable as the embedding inferred by Mercator, the majority of plausible embeddings are less reliable.
@@
The hierarchical organization of graphs is another example of complex properties captured by hyperbolic embeddings. The global hierarchy level was introduced to quantify this, using a function of the angular distance between neighboring vertices with coordinates in the outer and inner parts of the embedding [77]. Figure 5g illustrates how, for the synthetic graph analyzed, most plausible embeddings have smaller global hierarchy levels than both the ground truth and the Mercator embeddings.
Finally, embeddings for the $\mathbb{S}^{1}$ model can be viewed as classifiers for edges since the likelihood maps potential edges (binary classes) to edge probabilities. If the embedding is a good representation of the graph, edges should have a high probability, while non-edges should have a low probability. We test the performance in the edge prediction task by generating many embeddings and synthetic graphs of 30 vertices and removing 5% of its edges. We then compute the rank of the removed edges among all the unconnected pairs, using their inferred connection probability, Eq. (1). If the removed edges still have a relative high existence probability, the embedding is a good predictor and the ranks should approach $1$. Figure 5h shows that Mercator and BIGUE perform almost identically and, unsurprisingly, the original embedding performs better. This suggests that both embeddings are robust to small perturbations.
III.2 Multimodality